OpenStudy (anonymous):

3. Boeing plans to increase its prices for Jetliners. With a selling price of $201.5 million and a cost of $190.1 million, what was the percent markup based on cost? Round to the nearest percent?

OpenStudy (anonymous):

What does "markup" mean in relation to a selling price vs a cost of production?

OpenStudy (anonymous):

I'm assuming it costs $190.1 and they sell it at $201.5 so the question is asking what is the percent it was marked up from the $190.1 to the $201.5

OpenStudy (anonymous):

exactly. if I have 2 numbers, 10 and 5, 10 is what percent of five?

OpenStudy (anonymous):

I have no clue I suck at math.

OpenStudy (anonymous):

:) 10 is 200% of 5 - 5/5 = 1 -> 1*100 = 100. This shows 5 is 100% of five. Makes sense. 10/5 = 2 -> 2*100 = 200. This shows 10 is 200% of five. If I make something for 5 dollars, and sell it for 10, my percent markup is 200% - 100% = 100% If I make something for 5 dollars and sell it for 8, my percent markup is (8/5)*100 - 100% = 60% Does that make sense?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So if I make something for $190.1 and sell it for $201.5.....

OpenStudy (anonymous):

So do I divide the 201.5 by 190.1 ? then times it by 100 then subtract 100?

OpenStudy (anonymous):

exactly. If you want to "save time," you can do it like \[ \\ 100 \left ( \frac{201.5}{190.1} -1 \right) \] since \[ 100 \left( \frac{201.5}{190.1} \right) - 100 = 100 \left ( \frac{201.5}{190.1} -1 \right) \] But that's correct!

OpenStudy (anonymous):

so what is the final answer rounded to the nearest percent?

OpenStudy (anonymous):

\[ 1901 \ ) \overline{2015} \]

OpenStudy (anonymous):

-1 times 100

OpenStudy (anonymous):

5% ?

OpenStudy (anonymous):

6& ; you'd round up since it's 5.997 :)

OpenStudy (anonymous):

~ 6%